# An Efficient Sparse Kernel Generator for O(3)-Equivariant Deep Networks

### v3

@@ Abstract
Rotation equivariant graph neural networks, i.e. networks designed to guarantee certain geometric relations between their inputs and outputs, yield state of the art performance on spatial deep learning tasks. They exhibit high data efficiency during training and significantly reduced inference time for interatomic potential calculations compared to classical approaches. Key to these models is the Clebsch-Gordon (CG) tensor product, a kernel that contracts two dense feature vectors with a highly-structured sparse tensor to produce a dense output vector. The operation, which may be repeated millions of times for typical equivariant models, is a costly and inefficient bottleneck. We introduce a GPU sparse kernel generator for the CG tensor product that provides significant speedups over the best existing open and closed-source implementations. Our implementation achieves high performance by carefully managing the limited GPU shared memory through static analysis at model compile-time, minimizing reads and writes to global memory. We break the tensor product into a series of smaller kernels with operands that fit entirely into registers, enabling us to emit long arithmetic instruction streams that maximize instruction-level parallelism. By fusing the CG tensor product with a subsequent graph convolution , we reduce both intermediate storage and global memory traffic over naïve approaches that duplicate input data. We also provide optimized kernels for the gradient of the CG tensor product and a novel identity for the higher partial derivatives required to predict interatomic forces. Our fused kernels offer up to 4.8x speedup for the forward pass and 3x for the backward pass over NVIDIA’s closed-source cuEquivariance package, as well as $>10$x speedup over the widely-used e3nn package. We offer up to 5.4x inference-time speedup for the MACE chemistry foundation model over the original unoptimized version.
@@ 1 Introduction
We introduce an open source CUDA kernel generator for the Clebsch-Gordon tensor product and its gradients. Compared to the popular e3nn package [10] that is widely used in equivariant deep learning models, we offer up to one order of magnitude improvement for both forward and backward passes. Our forward kernels also exhibit up to 2x speedup over NVIDIA’s closed-source cuEquivariance package [11] on configurations used in graph neural networks, improving to 4.5x when we fuse the tensor product with the subsequent message passing operation. Our key innovations include:
@@
Each nonzero block of $\mathcal{P}$ in Figure 1 is a structured sparse tensor (see Figure 2 for illustrations of the nonzero pattern). Popular codes [10, 11] fill these blocks with explicit zeros and use optimized dense linear algebra primitives to execute the tensor product, performing unnecessary work in the process. By contrast, we use Just-in-Time (JIT) compilation to generate kernels that only perform work for nonzero tensor entries, achieving significantly higher throughput as block sparsity increases (see Table 2). While previous works [19, 17] use similar fine-grained approaches to optimize kernels for each block in isolation, we achieve high throughput by JIT-compiling a single kernel for the entire sparse tensor $\mathcal{P}$. By compiling kernels optimized for an entire sequence of nonzero blocks, we exhibit a degree of instruction level parallelism and data reuse that prior approaches do not provide.
@@
We demonstrate benefits far beyond reduced kernel launch overhead by fusing the CG tensor product with a graph convolution kernel (a very common pattern [34, 4, 2]). We embed the CG tensor product and its backward pass into two algorithms for sparse-dense matrix multiplication: a simple, flexible implementation using atomic operations and a faster deterministic version using a fixup buffer. As a consequence, we achieve significant data reuse at the L2 cache level and a reduction in global memory writes.
@@ 2.1 Representations of O(3)
Figure 2: Three examples of CG coefficient tensors depicted by blue cubes in Figure 1. The fraction of zero entries increases with the tensor order, and the full CG tensor contains several copies of the blocks pictured here.
@@ 2.5 Context and Related Work
The cuEquivariance package [11], released by NVIDIA concurrent to this project’s development, offers the fastest implementation of the CG tensor product outside of our work. However, their kernels are closed-source and do not appear to exploit sparsity within each nonzero block of $\mathcal{P}$. Our kernel performance matches or exceeds cuEquivariance, often by substantial margins. Other relevant codes include GELib [19], Sphericart [5], and Equitriton [23]. The former efficiently computes CG tensor products, while the latter two accelerate spherical harmonic polynomial evaluation. Although GELib is less widely used in production models, we analyze its custom CUDA kernels in Section 3.6.
@@
The intense cost of the CG tensor product has fueled the search for cheap yet accurate algorithms. For example, Schütt et al. [31] propose an equivariant message passing neural network that operates in Cartesian space, eliminating the need for CG tensor products. For SO(3)-equivariant graph convolution, Passaro and Zitnick [28] align the node embeddings with spherical harmonic edge features before interacting the two. This innovation sparsifies the interaction tensor and asymptotically decreases the cost of each tensor product. Luo et al. [26] also produce asymptotic speedups by connecting the tensor product with a spherical harmonic product accelerated through the fast Fourier transform, an operation they call the Gaunt tensor product. Xie et al. [38] counter that the Gaunt tensor product does not produce results directly comparable to the CG tensor product, and that the former may sacrifice model expressivity. Lee et al. [22] observe that higher-order equivariant features may not improve model performance, while Qu and Krishnapriyan [30] achieve high accuracy by eliminating hard rotational constraints in atomic potential models.
@@
GPUs execute kernels by launching a large number of parallel threads running the same program, each accessing a small set of local registers. In a typical program, threads load data from global memory, perform computation with data in their registers, and store back results. Kernels are most efficient when groups of 32-64 threads (called “warps”) execute the same instruction or perform memory transaction on contiguous, aligned segments of data. Warps execute asynchronously and are grouped into cooperative thread arrays (CTAs, called blocks in CUDA), which can communicate through a fast, limited pool of shared memory. Warps can synchronize at the CTA level, but the synchronization incurs overhead.
@@ 3.1 Computation Scheduling
    Otherwise, we use a greedy algorithm. In each phase, the shared memory pool is filled with as many segments of $\bm{x}$, $\bm{y}$, $\bm{W}$ and $\bm{z}$ that can fit. Between phases, segments are flushed and reloaded as needed.
@@ 3.2 JIT Subkernel Implementations
The matrix multiplication by the weights at the end of Algorithm 2 depends on the structure of $\bm{W}$. When $\bm{W}$ is square and diagonal (kernel B), multiplication proceeds asynchronously in parallel across all threads. When $\bm{W}$ is a general dense matrix, we temporarily store $\bm{z}_{\textrm{reg}}$ to shared memory and perform a warp-level matrix-multiplication across all threads. Our final implementation uses a mix of hand-rolled kernels and CUTLASS [33] for warp matrix multiplication.
@@ 3.4 Higher Partial Derivatives
Our analysis so far has focused on the forward and backward passes for the CG tensor product. For interatomic potential models, we require higher-order derivatives to optimize force predictions during training [16], as we explain below. Rather than write new kernels for these derivatives, we provide a novel (to the best of our knowledge) calculation that implements them using the existing forward and backward pass kernels.
@@
Our approach offers several additional advantages. Equations (3.4) and (3.5) are agnostic to the structure of $\mathcal{P}$, rendering the double-backward algorithm correct as long as the associated forward and backward kernels are correct. Because Equation (3.4) recursively calls operations with registered autograd formulas, an autograd framework can compute arbitrary higher partial derivatives of our model, not just double-backward. Finally, these formulas continue to apply when the CG tensor product is embedded within a graph convolution (see Section 3.5), composing seamlessly with the other optimizations we introduce.
@@ 3.5 Graph Convolution and Kernel Fusion
The second algorithm, which we omit for brevity, functions almost identically to Algorithm 4, but replaces the fixup / store logic with an atomic accumulation at every inner loop iteration. This method performs $O(\left|E\right|)$ atomic storebacks, but does not require a sorted input graph or adjacency transpose permutation. Figure 8 shows that the atomic algorithm still benefits from increased temporal L2 cache locality induced by the graph.
3.6 Analysis and Related Work Comparison
Our JIT-based approach embeds the sparse tensor structure and values into the GPU instruction stream, which is only possible because $\mathcal{P}$ has relatively few nonzero entries in each block. Because GPU instructions must also be fetched from global memory, performance eventually degrades as blocks exceed several thousand nonzero entries. We do not encounter such tensors in practical models.
Kondor et al. [20] used the GElib library [19] to implement one of the original $\textrm{O}(3)$-equivariant deep learning models. Their library also employs fine-grained loop unrolling in CUDA and parallelizes distinct tensor product evaluations across threads. Despite these strong innovations, their statically compiled library does not provide an optimized backward pass, synchronizes at the CTA rather than warp level, and handles each nonzero block of $\mathcal{P}$ with a separate kernel invocation. GElib does not match the performance of e3nn and faces obstacles to data reuse, which we overcome through our computation schedule and JIT compilation.
@@ 4 Experiments
Our kernel generator is available online111 https://github.com/PASSIONLab/OpenEquivariance as an installable Python package. We adopted the frontend interface of e3nn [9, 10] and used QuTiP [14, 21] to generate CG coefficients; we thank the authors of both packages. We tested correctness against e3nn to ensure that our kernels produce identical results, up to floating point roundoff and a well-defined permutation of the weights $\bm{W}$ on certain input configurations. In cases where weight reordering is required, we provide a function for easy migration. We use the NVIDIA Runtime Compiler (NVRTC) to compile our generated kernels through a C++ extension to Python.
Experiments were conducted on NVIDIA A100 GPU nodes of NERSC Perlmutter (each equipped with an AMD EPYC 7763 CPU). Table 1 lists the advertised maximum memory bandwidth and compute peaks for multiple datatypes, a yardstick for our results.
@@
As baselines, we used the PyTorch versions of e3nn (v0.5.4) [10] and NVIDIA cuEquivariance (v0.0.2) [11]. The e3nn implementation was accelerated with torch.compile except where prohibited by memory constraints. For Figures 5 and 6, we benchmarked all functions through a uniform PyTorch interface and included any overhead in the measured runtime. Figures 7, Table 2, and Figure 9 (right) rely on kernel runtime measurements without PyTorch overhead.
4.1 Throughput Comparison
Figure 5: Throughput of CG tensor products (batch size 50K), kernel B configurations without SpMM kernel fusion. On difficult configurations like Nequip-benzene with massive output vector lengths, we exhibit up to 10x improvement over e3nn and cuE.
@@ 4.1 Throughput Comparison
We first profiled our kernels on a large collection of model configurations used by Nequip [4] and MACE [2]. For each model, we selected the most expensive tensor product to benchmark.
Figure 5 shows the results of our profiling on configurations that use only Kernel B (see Figure 3). Our median FP32 speedup over e3nn was 5.7x, resp. 5.0x, for the forward and backward passes, with a maximum of 9.3x for the forward pass. We observed a median speedup of 1.7x over cuE for the FP32 forward pass, which drops to 1.2x in FP64 precision. We achieve parity with cuE on the larger MACE model in FP64 precision; here, both cuE and our kernels have nearly identical runtime. We explore this configuration further in our kernel fusion benchmarks. Our implementation exhibits an outsized speedup (10-20x) on the Nequip-benzene configuration over both baselines. Here, the large output vector size ($\sim 20\textrm{k}$ data words) requires careful shared memory management, which our scheduling algorithm provides.
To benchmark kernel C, we used the Tetris polynomial from e3nn’s documentation [10] and two configurations based on DiffDock [7]. We exhibit between 1.4x and 2.1x speedup over cuE for both forward and backward passes on DiffDock. The Tetris polynomial uses an identical weight matrix $\bm{W}$ for all inputs $\bm{x}_{1}...\bm{x}_{B}$ and $\bm{y}_{1}...\bm{y}_{B}$, a feature unique to this model. We use atomic accumulation in the backward pass to compute the gradient with respect to the shared weights, which reduces our relative speedup in FP64 precision. Our speedups over all configurations are less dramatic for kernel C, which has a workload dominated by the small dense matrix multiplication in Algorithms 2 and 3.
@@ 4.2 Roofline Analysis
ID Description Dir. AI TFLOP/s
cuE ours
[…]
B 1.2 0.43 1.84
[…]
B 2.1 0.73 3.22
[…]
B 4.1 1.19 6.22
[…]
B 7.4 1.39 9.77
[…]
B 6.5 1.35 8.94
[…]
B 6.9 1.41 9.24
[…]
B 9.5 1.46 10.67
[…]
B 12.3 1.56 11.30
Table 2: Arithmetic intensity (AI) and throughput for our kernels vs. cuEquivariance (cuE) for varied problem configurations, FP32 forward and backward passes.
@@
Figure 7 and Table 2 show our profiling results, which indicate consistently high compute and bandwidth utilization for our kernels. In the bandwidth-limited portion of the diagram, our implementation increases its throughput linearly to match increasing arithmetic intensity in the input configurations. By contrast, the throughput of cuEquivariance’s forward pass stagnates. Our kernel performance saturates at 58% of the FP32 peak, possibly because Algorithms 2 and 3 contain contain a significant fraction of non fused-multiply-add (FMA) instructions.
Figure 7: Roofline analysis of our kernels on the input configurations from Table 2, batch size 200k. Our kernel performance closely tracks the slope of the global memory roofline, whereas cuE plateaus early in the memory-bandwidth limited region of the plot.
@@ 4.3 Kernel Fusion Benchmarks
Figure 8: Speedup of our fused MACE convolution kernels over cuE-scattersum, which calls cuEquivariance and follows it by a scatter-sum operation.
@@
Figure 8 shows the speedup of our fused implementations benchmarked on the most expensive tensor product in the MACE-large model. Our baseline, “cuE-scattersum”, implements the unfused strategy in Section 3.5 by duplicating node embeddings, executing a large batch of tensor products with cuEquivariance, and finally performing row-based reduction by keys. The “fast-scattersum” implementation adopts the same method with our kernels instead of our cuEquivariance. Without kernel fusion, our implementation and cuEquivariance exhibit nearly identical performance across all configurations except the backwards pass in FP32 precision (we are 2x faster). The performance gap is smaller than in Figure 5 since the scatter-sum accounts for 50-60% of the runtime.
The“fast-fused-det” and “fast-fused-atomic” implementations use the memory-efficient formulations from Section 3.5, offering dramatically faster performance as a result. Our deterministic fused algorithm offers up to 4.8x speedup over cuEquivariance for the forward pass and 3.5x on the backward pass. Our poorest performance occurs with the atomic algorithm in FP64 precision for the backward pass, where we still see 1.3x-1.4x speedup over cuEquivariance.
4.4 Acceleration of MACE
The MACE foundation model [3], trained on data from the Materials Project [13], implements the equivariant graph neural network architecture depicted in Figure 4B. We patched MACE to sort nonzeros of its atomic adjacency matrix according to compressed sparse row (CSR) order, as well as compute the transpose permutation. We then substituted Algorithm 4 in place of the existing graph convolution to measure performance relative to other kernel providers. Our benchmark was conducted on the carbon lattice listed in Table 3.

### v4

@@ Abstract
Rotation equivariant graph neural networks, i.e. networks designed to guarantee certain geometric relations between their inputs and outputs, yield state of the art performance on spatial deep learning tasks. They exhibit high data efficiency during training and significantly reduced inference time for interatomic potential calculations compared to classical approaches. Key to these models is the Clebsch-Gordon (CG) tensor product, a kernel that contracts two dense feature vectors with a highly-structured sparse tensor to produce a dense output vector. The operation, which may be repeated millions of times for typical equivariant models, is a costly and inefficient bottleneck. We introduce a GPU sparse kernel generator for the CG tensor product that provides significant speedups over the best existing open and closed-source implementations. Our implementation achieves high performance by carefully managing the limited GPU shared memory through static analysis at model compile-time, minimizing reads and writes to global memory. We break the tensor product into a series of smaller kernels with operands that fit entirely into registers, enabling us to emit long arithmetic instruction streams that maximize instruction-level parallelism. By fusing the CG tensor product with a subsequent graph convolution , we reduce both intermediate storage and global memory traffic over naïve approaches that duplicate input data. We also provide optimized kernels for the gradient of the CG tensor product and a novel identity for the higher partial derivatives required to predict interatomic forces. Our kernels offer up to 1.3x speedup for the over NVIDIA’s closed-source cuEquivariance package, as well as $>10$x speedup over the widely-used e3nn package. In FP64 precision, we offer up to 6.2x inference-time speedup for the MACE chemistry foundation model over the original unoptimized version.
@@ 1 Introduction
We introduce an open source kernel generator for the Clebsch-Gordon tensor product on both NVIDIA and AMD GPUs. Compared to the popular e3nn package [10] that is widely used in equivariant deep learning models, we offer up to one order of magnitude improvement for both forward and backward passes. Our kernels also exhibit up to 1.0-1.3x speedup over NVIDIA’s closed-source cuEquivariance v0.4.0 package [11] on configurations used in graph neural networks, although our second derivative kernel is slower by 30% on certain inputs. Our key innovations include:
@@
Each nonzero block of $\mathcal{P}$ in Figure 1 is a structured sparse tensor (see Figure 2 for illustrations of the nonzero pattern). Popular existing codes [10] fill these blocks with explicit zeros and use optimized dense linear algebra primitives to execute the tensor product, performing unnecessary work in the process. By contrast, we use Just-in-Time (JIT) compilation to generate kernels that only perform work for nonzero tensor entries, achieving significantly higher throughput as block sparsity increases. While previous works [19, 17] use similar fine-grained approaches to optimize kernels for each block in isolation, we achieve high throughput by JIT-compiling a single kernel for the entire sparse tensor $\mathcal{P}$. By compiling kernels optimized for an entire sequence of nonzero blocks, we exhibit a degree of instruction level parallelism and data reuse that prior approaches do not provide.
@@
We demonstrate benefits far beyond reduced kernel launch overhead by fusing the CG tensor product with a graph convolution kernel (a very common pattern [32, 4, 2]). We embed the CG tensor product and its backward pass into two algorithms for sparse-dense matrix multiplication: a simple, flexible implementation using atomic operations and a faster deterministic version using a fixup buffer. As a consequence, our work is the first to reap significant model memory savings, a reduction in global memory writes, and data reuse at the L2 cache level.
@@ 2.1 Representations of O(3)
Figure 2: Three examples of coefficient tensors depicted by blue cubes in Figure 1. The fraction of zero entries increases with the tensor order, and the full CG tensor contains several copies of the blocks pictured here.
@@ 2.5 Context and Related Work
The cuEquivariance package [11], released by NVIDIA concurrent to this project’s development, offers the fastest implementation of the CG tensor product outside of our work. However, their kernels are closed-source and do not appear to exploit sparsity within each nonzero block of $\mathcal{P}$. Our kernel performance matches or exceeds cuEquivariance, often by substantial margins. Other relevant codes include GELib [19], Sphericart [5], and Equitriton [22]. The former efficiently computes CG tensor products, while the latter two accelerate spherical harmonic polynomial evaluation.
@@
The intense cost of the CG tensor product has fueled the search for cheap yet accurate algorithms. For example, Schütt et al. [29] propose an equivariant message passing neural network that operates in Cartesian space, eliminating the need for CG tensor products. For SO(3)-equivariant graph convolution, Passaro and Zitnick [27] align the node embeddings with spherical harmonic edge features before interacting the two. This innovation sparsifies the interaction tensor and asymptotically decreases the cost of each tensor product. Luo et al. [25] also produce asymptotic speedups by connecting the tensor product with a spherical harmonic product accelerated through the fast Fourier transform, an operation they call the Gaunt tensor product. Xie et al. [36] counter that the Gaunt tensor product does not produce results directly comparable to the CG tensor product, and that the former may sacrifice model expressivity.
@@
GPUs execute kernels by launching a large number of parallel threads running the same program, each accessing a small set of local registers. In a typical program, threads load data from global memory, perform computation with data in their registers, and store back results. Kernels are most efficient when groups of 32-64 threads (called “warps”) execute the same instruction or perform memory transaction on contiguous, aligned segments of data. Warps execute asynchronously and are grouped into cooperative thread arrays (CTAs) that communicate through a fast, limited pool of shared memory. Warps can synchronize at the CTA level, but the synchronization incurs overhead.
@@ 3.1 Computation Scheduling
    Otherwise, we use a greedy algorithm. In each phase, the shared memory pool is filled with as many segments of $\bm{x}$, $\bm{y}$, $\bm{W}$ and $\bm{z}$ that can fit. Segments are flushed and reloaded as needed.
@@ 3.2 JIT Subkernel Implementations
The matrix multiplication by the weights at the end of Algorithm 2 depends on the structure of $\bm{W}$. When $\bm{W}$ is square and diagonal (kernel B), multiplication proceeds asynchronously in parallel across all threads. When $\bm{W}$ is a general dense matrix, we temporarily store $\bm{z}_{\textrm{reg}}$ to shared memory and perform a warp-level matrix-multiplication across all threads.
@@ 3.4 Higher Partial Derivatives
For interatomic potential models, we require higher-order derivatives to optimize force predictions during training [16], as we explain below. Rather than write new kernels for these derivatives, we provide a novel (to the best of our knowledge) calculation that implements them using the existing forward and backward pass kernels.
@@ 3.5 Graph Convolution and Kernel Fusion
The second algorithm, which we omit for brevity, functions almost identically to Algorithm 4, but replaces the fixup / store logic with an atomic accumulation at every inner loop iteration. This nondeterministic method performs $O(\left|E\right|)$ atomic storebacks, but does not require a sorted input graph or adjacency transpose permutation.
@@ 4 Experiments
Our kernel generator is available online111 https://github.com/PASSIONLab/OpenEquivariance as an installable Python package. We adopted the frontend interface of e3nn [9, 10] and used QuTiP [13, 21] to generate CG coefficients. We tested correctness against e3nn to ensure that our kernels produce identical results, up to floating point roundoff and a well-defined permutation of the weights $\bm{W}$ on certain input configurations. In cases where weight reordering is required, we provide a function for easy migration. We use the NVIDIA and AMD HIP Runtime Compilers to compile our generated kernels through a C++ extension to Python.
@@
The majority of experiments were conducted on NVIDIA A100 GPU nodes of NERSC Perlmutter (each equipped with an AMD EPYC 7763 CPU). Table 1 lists the advertised maximum memory bandwidth and compute peaks for multiple datatypes, a yardstick for our results. Section 4.4 covers performance on other GPU models.
As baselines, we used the PyTorch versions of e3nn (v0.5.6) [10] and NVIDIA cuEquivariance (v0.4.0) [11]. The e3nn implementation was accelerated with torch.compile except where prohibited by memory constraints. For Figures 5, 6, 7, and 9, we benchmarked all functions through a uniform PyTorch interface and included any overhead in the measured runtime. Figures 8 and 10 (right) rely on kernel runtime measurements without PyTorch overhead.
cuEquivariance experienced a significant efficiency increase since v0.2.0, the latest version available when our first preprint was released (see Figures 8, 10). Since that early release, the authors also added JIT capability and fused convolution, although the closed source kernel backend renders the details opaque. Unless otherwise noted, we report all benchmarks against cuE v0.4.0.
4.1 Forward / Backward Throughput
Figure 5: Throughput of CG tensor products (batch size 50K), kernel B configurations without SpMM kernel fusion. On difficult configurations like Nequip-benzene with massive output vector lengths, we exhibit more than 10x improvement over e3nn.
@@ 4.1 Forward / Backward Throughput
We first profiled our kernels on a large collection of model configurations used by Nequip [4] and MACE [2]. For each model, we selected an expensive representative tensor product to benchmark. Figure 5 shows the results of our profiling on configurations that use only Kernel B (see Figure 3). Our median FP32 speedup over e3nn was 5.9x, resp. 4.8x, for the forward and backward passes, with a maximum of 9.2x for the forward pass. We observed a median speedup of 1.6x over cuE for the FP32 forward pass, which drops to 1.3x in FP64 precision. Our median performance approaches parity with cuE on the backward pass, with a minimum and maximum speedup of 0.72x and 1.32x in FP64 precision.
To benchmark kernel C, we used the Tetris polynomial from e3nn’s documentation [10] and two configurations based on DiffDock [7]. We exhibit between 1.4x and 2.0x speedup over cuE for both forward and backward passes on DiffDock. Our speedups over e3nn are less dramatic for kernel C, which has a workload dominated by the small dense matrix multiplication in Algorithms 2 and 3.
4.2 Second Derivative Performance
We analyzed the double backward pass for the tensor products from the prior section (excluding the Tetris polynomial, which does not require it). Figure 7 shows our results. Across datatypes and model configurations, our speedup ranges from 5.5x to 35x over e3nn and 0.69x-1.69x over cuE. Although our median speedup over cuE is 0.73x in FP32 precision and 0.93x for FP64 precision, we exhibit lower runtime on all DiffDock tensor products and several Nequip configurations.
For many Nequip / MACE configurations, the performance gap between our implementation and cuE could likely shrink with some judicious kernel tuning. In particular, we could improve our heuristic selection of the warp count per block, the number of blocks, and the shared memory allotted to each block. We leave tuning these hyperparameters as future work.
Figure 7: Throughput of second derivative kernels (batch size 20K) for chemistry / protein models.
@@ 4.3 Roofline Analysis
Figure 8 shows our results, which indicate consistently high compute and bandwidth utilization for both our kernels and the latest version of cuE. An earlier version of cuE (v0.2.0, indicated on the plot as cuE-old) exhibited significantly lower efficiency, which has since been corrected by the package authors. The performance of all kernels saturates at 58% of the FP32 peak, likely because Algorithms 2 and 3 contain contain a significant fraction of non fused-multiply-add (FMA) instructions.
Figure 8: Roofline analysis for input configurations of varying arithmetic intensity, batch size 200K. Our kernels and the latest version of cuE closely track the slope of the global memory roofline, indicating high efficiency. An older version of cuE (v0.2.0) is also included to highlight the performance improvement in their package (see top of Section 4).
4.4 Additional GPU Models
We also tested our kernels on the NVIDIA A5000 and a single GPU die of the AMD MI250x. Table 2 compares the MACE tensor product runtime across architectures and kernel providers; our codebase contains a more complete set of benchmarks. The A5000 performance matches our expectations given its lower memory bandwidth compared to the A100. While we also saw significant speedup on the MI250x, we detected somewhat lower memory bandwidth utilization than predicted.
GPU forward backward
e3nn cuE ours e3nn cuE ours
A100 13 2.8 2.0 21 3.5 3.7
A5000 29 4.2 3.8 42 9.3 11
MI250x 41 - 3.0 128 - 15
Table 2: MACE-large isolated tensor product runtime (ms), batch size 50K, FP32 unfused.
@@ 4.5 Kernel Fusion Benchmarks
Figure 9: Speedup of convolution kernels over cuE-scattersum, which calls cuEquivariance and follows it by a scatter-sum operation. cuE-fused refers to a new kernel introduced in cuE v0.4.0.
@@
Figure 9 shows the speedup of fused implementations benchmarked on the most expensive tensor product in the MACE-large model. The baseline, “cuE-scattersum”, implements the unfused strategy in Section 3.5 by duplicating node embeddings, executing a large batch of tensor products with cuEquivariance, and finally performing row-based reduction by keys. Our deterministic fused algorithm offers the greatest speedup in FP64 precision on the carbon lattice forward pass over cuE (roughly 1.3x speedup). On the other hand, cuE-fused offers 1.3-1.4x speedup on the backward pass, and we aim to close this performance gap.
4.6 Acceleration of Nequip and MACE
The Nequip [4, 31] and MACE [2, 3] interatomic potential models implement the equivariant graph neural network architecture in Figure 4B. Both have similar message passing structures, while MACE incorporates higher order interactions for its node features. Our first benchmark uses the Nequip-ASE calculator interface to evaluate forces on a large box of water molecules. Due to recent updates to Nequip’s software [31] and time constraints, we only used the nondeterministic fused convolution for our measurements, which appear in Table 4. For simplicity, we report speedup with respect to the Nequip Python interface without JITScript or torch.compile, although our package is fully compatible with these subsystems.
GPU Speedup over Unmodified Nequip
ours-unfused ours-fused
A100 6.3x 7.8x
MI250x 3.9x 4.4x
Table 4: Force evaluation speedup for our kernels on a 4-layer FP32 Nequip model, 5184-atom water box system.
For MACE, we patched the code to sort nonzeros of the atomic adjacency matrix according to compressed sparse row (CSR) order. We then substituted our deterministic fused convolution into the model and conducted our benchmark on the carbon lattice in Table 3. MACE uses a distinct set of weight matrices for each atomic species, and an inefficiency in the baseline code causes its runtime to increase disproportionately to the useful computation involved. Our model has a species dictionary of eight elements (to trigger the problem in the baseline code, even though the carbon lattice only requires one), and our package includes a module to optimize away the inefficiency.